OpenStudy (anonymous):

decide whether the pair of lines is parallel, perpendicular or neither 2x+4y=6 2x+4y=1 I said they are parallel

OpenStudy (anonymous):

yes same slopes

OpenStudy (anonymous):

determine the slope. Parallel lines have the same slope. y=-1/2x+3/2 y=-1/2x+1/4 so you are correct

OpenStudy (anonymous):

2x+4y=6--->y=-1/2x+3/2 2x+4y=1--->y=-1/2x+1/4 they have same gradient that is m so these lines are parallel

OpenStudy (anonymous):

In general, the line ax + by + c = 0 has slope of - a/b. Then the slope of your first line is - 2/4 or -1/2 and so is the second therefore the lines are parallel. A line perpendicular to another with slope m would have a slope of - 1/m (ie the product of the slopes is -1).